John left school for home going 10km/h on his bike. Steve left 20 minutes later going 20km/h in order to catch up to John. How long did it take fore Steve to catch up to John
Answer by sachi(548)


let after t min john starts Steve catch up John
in t min
John travels the distance =(10/60)*t km
& Steve travels the distance =(20/60)*(t-20)km
as they meet the distance traveled by both are same
so (10/60)*t km=(20/60)*(t-20)km
or t=2(t-20)=2t-40
or 2t-t=40
ot t=40 min
so after 40 min of John starts Or 20 min of Steve starts they meet each other